# SINET: Sparsity-driven Interpretable Neural Network for Underwater Image Enhancement

### v1

@@ Abstract
Improving the quality of underwater images is essential for advancing marine research and technology. This work introduces a sparsity-driven interpretable neural network (SINET) for the underwater image enhancement (UIE) task. Unlike pure deep learning methods, our network architecture is based on a novel channel-specific convolutional sparse coding (CCSC) model, ensuring good interpretability of the underlying image enhancement process. The key feature of SINET is that it estimates the salient features from the three color channels using three sparse feature estimation blocks (SFEBs). The architecture of SFEB is designed by unrolling an iterative algorithm for solving the $\ell_{1}$ regulaized convolutional sparse coding (CSC) problem. Our experiments show that SINET surpasses state-of-the-art PSNR value by $1.05$ dB with $3873$ times lower computational complexity.
@@ I Introduction
where $D(\cdot)$ is a convolutional dictionary which can be learned and $K$ is the number of convolution filters. $H$ and $W$ are the image height and width, respectively. In Eqn. 1, $z$ is enforced to be sparse by $\ell_{0}$ or $\ell_{1}$ norm regularization, which helps to highlight the salient features in the input image $I$. Optimization methods like the alternating direction method of multipliers (ADMM) [10] are prevailing approaches for solving Eqn. 1. However, such algorithms are time-consuming, limiting their large-scale data applications.
To incorporate the advantages of optimization and deep learning methods, algorithm unrolling-based methods have been developed [11, 12]. These methods unroll the iteration steps of an optimization problem into a deep neural network. CSC-based algorithm unrolling networks have been applied in several image processing tasks like image denoising [13] and super-resolution [14]. Since the CSC-based unrolling network provides an interpretable and efficient framework, it can also be effective for UIE. However, for UIE, addressing the unique challenge of wavelength-dependent light attenuation is important. In this context, employing a separate CSC model for each color channel can be highly effective.
@@ II-A Problem Formulation
In Eqn. 2, we constrain $z_{i}$ to be $\ell_{1}$ regularized. The $\ell_{1}$ regularization is a widely used technique that promotes sparsity in the features and thus highlights the salient features. So, we estimate $z_{i}$ by solving the following Eqn.,
@@
Eqn. 4 is an $\ell_{1}$ regularized CSC problem, for solving which we propose a sparse feature estimation block described in the next subsection.
@@ II-B Sparse Feature Estimation Block (SFEB)
In SFEB, we estimate the sparse features by solving Eqn. 4. Sreter et al. [15] proposed a convolutional extension of the learned iterative shrinkage thresholding algorithm (LISTA) [16] to solve the $\ell_{1}$ regularized CSC problem. From [15], the iteration step for updating $z_{i}$ is,
$z_{i}^{k+1}=S_{\theta_{i}}\Big{(}z_{i}^{k}-W_{u_{i}}\Big{(}W_{d_{i}}\Big{(}z_{i}^{k}\Big{)}\Big{)}-W_{u_{i}}\Big{(}I_{s_{i}}\Big{)}\Big{)}$ (5)
@@
$sgn(\cdot)$ is the sign function. $\theta_{i}$ is the threshold value, which is a learnable parameter. As shown in Eqn. 6, $S_{\theta_{i}}(\cdot)$ nullifies the input values with an absolute values lower than $\theta_{i}$. Thus, it promotes sparsity and highlights the salient features in the input color channel.
However, in Eqn. 5, the learnable parameter $\theta_{i}$ and convolution layers $W_{u_{i}}(\cdot)$, $W_{d_{i}}(\cdot)$ are shared across all the iterations in [15]. Such parameter sharing limits the performance of the sparse estimation. Motivated by [17], we use different $\theta_{i}$ and different convolution layers for each iteration to improve the estimation accuracy. Additionally, for the input $I_{s_{i}}$, using a learnable layer that is different from $W_{u_{i}}(\cdot)$ can have better performance [18]. Incorporating these modifications in Eqn. 5, the iteration step for updating $z_{i}$ becomes,
$z_{i}^{k+1}=S_{\theta_{i}^{k}}\Big{(}z_{i}^{k}-W_{u_{i}}^{k}\Big{(}W_{d_{i}}^{k}\Big{(}z_{i}^{k}\Big{)}\Big{)}-W_{in_{i}}^{k}\Big{(}I_{s_{i}}\Big{)}\Big{)}$ (7)
@@
where $sp(\cdot)$ is the softplus function, $w_{\theta_{i}}$ and $b_{\theta_{i}}$ are learnable parameters. We unfold Eqn. 7 into the iteration module (IM) shown in Fig. 1. Assuming $z_{i}^{k}=0;\>\>\forall k<0$ and stacking a number of iteration modules, we construct our sparse feature estimation block (SFEB).
@@ III-A Setup
We utilize publicly available datasets for training and testing SINET. For training, we use 890 images from the UIEB [1] dataset. We train SINET using Adam optimizer with a constant learning rate of $1\times 10^{-4}$ and batch size of $4$. While training, the images are randomly cropped to the resolution of $256\times 256$. For image augmentation, we utilize horizontal and vertical flipping. We test the performance on 60 images from the LSUI [20] dataset without finetuning to check the model’s generalization ability. Moreover, to check the model’s performance on real-world images, we test on 60 images from the UIEB-challenge (UIEBC) dataset, which has no ground truth images. We compare the PSNR and SSIM metrics for the LSUI dataset, which has ground truth images. For the UIEBC dataset, we compare the performance using the UIQM metric [21]. For the implementation of SINET, we set the kernel size for each convolutional layer as $11\times 1$ 11 and the number of filters as $K=16$. The number of iteration modules in SFEB is set to $4$. All experiments are conducted within the Pytorch framework using an NVIDIA A40 GPU.
@@ III-B Performance Comparison
We compare SINET with five state-of-the-art methods: UIEWD [2], SCNet [3], CLNet [22], DWNet [5], and XCNET [6]. Figure 2 shows the visual comparison results on two images from the LSUI dataset and one image from the UIEBC dataset. UIEWD produces severe color distortions. For the images from the LSUI dataset, SCNet, CLNet, DWNet, and XCNET change the greenish tone that is present in the GT image. Whereas SINET preserves the greenish tone and reconstructs the images with better structure similarity with the GT image. For the image from the UIEBC dataset, SCNet, DWNet, and XCNET change the green portion of the input image to a different color. Though CLNet preserves the green color, but it produces a blurry image. In contrast, SINET preserves the green color and produces an image with better structural details.

### v2

@@ Abstract
Improving the quality of underwater images is essential for advancing marine research and technology. This work introduces a sparsity-driven interpretable neural network (SINET) for the underwater image enhancement (UIE) task. Unlike pure deep learning methods, our network architecture is based on a novel channel-specific convolutional sparse coding (CCSC) model, ensuring good interpretability of the underlying image enhancement process. The key feature of SINET is that it estimates the salient features from the three color channels using three sparse feature estimation blocks (SFEBs). The architecture of SFEB is designed by unrolling an iterative algorithm for solving the $\ell_{1}$ regulaized convolutional sparse coding (CSC) problem. Our experiments show that SINET surpasses state-of-the-art PSNR value by $1.05$ dB with $3873$ times lower computational complexity. Code can be found at: https://github.com/gargi884/SINET-UIE/tree/main.
@@ I Introduction
where $D(\cdot)$ is a convolutional dictionary which can be learned and $K$ is the number of convolution filters. $H$ and $W$ are the image height and width, respectively. In Eqn. (1), $z$ is enforced to be sparse by $\ell_{0}$ or $\ell_{1}$ norm regularization, which helps to highlight the salient features in the input image $I$. Optimization methods like the alternating direction method of multipliers (ADMM) [10] are prevailing approaches for solving Eqn. (1). However, such algorithms are time-consuming, limiting their large-scale data applications.
To incorporate the advantages of optimization and deep learning methods, algorithm unrolling-based methods have been developed [11, 12]. These methods unroll the iteration steps of an optimization problem into a deep neural network. CSC-based algorithm unrolling networks have been applied in several image processing tasks like image denoising [13, 14] and super-resolution [15, 16]. Since the CSC-based unrolling network provides an interpretable and efficient framework, it can also be effective for UIE. However, for UIE, addressing the unique challenge of wavelength-dependent light attenuation is important. In this context, employing a separate CSC model for each color channel can be highly effective.
@@ II-A Problem Formulation
In Eqn. (2), we constrain $z_{i}$ to be $\ell_{1}$ regularized. The $\ell_{1}$ regularization is a widely used technique that promotes sparsity in the features and thus highlights the salient features. So, we estimate $z_{i}$ by solving the following Eqn.,
@@
Eqn. (4) is an $\ell_{1}$ regularized CSC problem, for solving which we propose a sparse feature estimation block described in the next subsection.
@@ II-B Sparse Feature Estimation Block (SFEB)
In SFEB, we estimate the sparse features by solving Eqn. (4). Sreter et al. [17] proposed a convolutional extension of the learned iterative shrinkage thresholding algorithm (LISTA) [18] to solve the $\ell_{1}$ regularized CSC problem. From [17], the iteration step for updating $z_{i}$ is,
$z_{i}^{k+1}=S_{\theta_{i}}\Big{(}z_{i}^{k}-W_{u_{i}}\Big{(}W_{d_{i}}\Big{(}z_{i}^{k}\Big{)}\Big{)}+W_{u_{i}}\Big{(}I_{s_{i}}\Big{)}\Big{)}$ (5)
@@
$sgn(\cdot)$ is the sign function. $\theta_{i}$ is the threshold value, which is a learnable parameter. As shown in Eqn. (6), $S_{\theta_{i}}(\cdot)$ nullifies the input values with an absolute values lower than $\theta_{i}$. Thus, it promotes sparsity and highlights the salient features in the input color channel.
However, in Eqn. (5), the learnable parameter $\theta_{i}$ and convolution layers $W_{u_{i}}(\cdot)$, $W_{d_{i}}(\cdot)$ are shared across all the iterations in [17]. Such parameter sharing limits the performance of the sparse estimation. Motivated by [19], we use different $\theta_{i}$ and different convolution layers for each iteration to improve the estimation accuracy. Additionally, for the input $I_{s_{i}}$, using a learnable layer that is different from $W_{u_{i}}(\cdot)$ can have better performance [20]. Incorporating these modifications in Eqn. (5), the iteration step for updating $z_{i}$ becomes,
$z_{i}^{k+1}=S_{\theta_{i}^{k}}\Big{(}z_{i}^{k}-W_{u_{i}}^{k}\Big{(}W_{d_{i}}^{k}\Big{(}z_{i}^{k}\Big{)}\Big{)}+W_{in_{i}}^{k}\Big{(}I_{s_{i}}\Big{)}\Big{)}$ (7)
@@
where $sp(\cdot)$ is the softplus function, $w_{\theta_{i}}$ and $b_{\theta_{i}}$ are learnable parameters. We unfold Eqn. (7) into the iteration module (IM) shown in Fig. 1. Assuming $z_{i}^{k}=0;\>\>\forall k<0$ and stacking a number of iteration modules, we construct our sparse feature estimation block (SFEB).
@@ III-A Setup
We utilize publicly available datasets for training and testing SINET. For training, we use 890 images from the UIEB [1] dataset. We train SINET using Adam optimizer with a constant learning rate of $1\times 10^{-4}$ and batch size of $4$. While training, the images are randomly cropped to the resolution of $256\times 256$. For image augmentation, we utilize horizontal and vertical flipping. We test the performance on 60 images from the LSUI [22] dataset without finetuning to check the model’s generalization ability. Moreover, to check the model’s performance on real-world images, we test on 60 images from the UIEB-challenge (UIEBC) dataset, which has no ground truth images. We compare the PSNR and SSIM metrics for the LSUI dataset, which has ground truth images. For the UIEBC dataset, we compare the performance using the UIQM metric [23]. For the implementation of SINET, we set the kernel size for each convolutional layer as $11\times 1$ 11 and the number of filters as $K=16$. The number of iteration modules in SFEB is set to $4$. All experiments are conducted within the Pytorch framework using an NVIDIA A40 GPU. A short demo of our work is available at: https://github.com/gargi884/SINET-UIE/blob/main/figs/demo.gif.
@@ III-B Performance Comparison
We compare SINET with five state-of-the-art methods: UIEWD [2], SCNet [3], CLNet [24], DWNet [5], and XCNET [6]. Figure 2 shows the visual comparison results on two images from the LSUI dataset and one image from the UIEBC dataset. UIEWD produces severe color distortions. For the images from the LSUI dataset, SCNet, CLNet, DWNet, and XCNET change the greenish tone that is present in the GT image. Whereas SINET preserves the greenish tone and reconstructs the images with better structure similarity with the GT image. For the image from the UIEBC dataset, SCNet, DWNet, and XCNET change the green portion of the input image to a different color. Though CLNet preserves the green color, but it produces a blurry image. In contrast, SINET preserves the green color and produces an image with better structural details. More visual comparisons can be found at: https://github.com/gargi884/SINET-UIE/blob/main/figs/visual.pdf.